OpenStudy (anonymous):

solve for x 2x^4+6x^3-50x^2-150x=0

OpenStudy (anonymous):

2x^3(x + 3) - 50x(x + 3) = 0 (x+3)(2x^3 -50x) = 0 (x+3)(2x(x^2 - 50) = 0 x = 0 or -3 also x^2 = 50 x = sqrt50 or -sqrt50

OpenStudy (anonymous):

summary zeros at x = 0 , -3 , 7.071, - 7.071.

OpenStudy (anonymous):

hold on araceli - theres an error on line 3 (x+3)(2x(x^2 - 25) = 0 its 25 not 50 so x^2 = 50 so x = 5 or -5 roots are 0, -3 , 5 and -5.

OpenStudy (anonymous):

correct this time

OpenStudy (anonymous):

sorry about that